calculate the apex of projectile relative to the ground if the initial vertical velocity of the object was 3m/s and the height of release was
Answer #1

Here

initial vertical velocity , v = 3 m/s

initial height = h0

apex of projectile relative to ground = h0 + v^2/(2 * g)

apex of projectile relative to ground = h0 + 3^2/(2 * 9.8)

apex of projectile relative to ground = h0 + 0.46